OpenStudy (anonymous):

How do i put this in factored form? 6x^2+4x

OpenStudy (anonymous):

the one way would be to use the quadratic formula.

OpenStudy (anonymous):

\[-b \pm \frac{ \sqrt{b^2-4ac} }{ 2a }\] a = 6 b = 4 c = 0

OpenStudy (anonymous):

okay cool thanks, I had that answer but I wasnt sure if i was supposed to have individual variables equal to certain numbers or if it was gonna look like (x+1)(x+4) (not those exact numbers but you get it lol)

OpenStudy (anonymous):

well x+1 is x = -1 so ya I get what you're saying :)